# EmojiDiff: Advanced Facial Expression Control with High Identity Preservation in Portrait Generation

### v2

@@ Abstract
This paper aims to bring fine-grained expression control while maintaining high-fidelity identity in portrait generation. This is challenging due to the mutual interference between expression and identity. On one hand, fine expression control signals inevitably introduce appearance-related semantics (e.g., facial contours, and ratio), which impact the identity of the generated portrait. On the other hand, even coarse-grained expression control can cause facial changes that compromise identity, since they all act on the face. These limitations remain unaddressed by previous generation methods, which primarily rely on coarse control signals or two-stage inference that integrates portrait animation. Here, we introduce EmojiDiff, the first end-to-end solution that enables simultaneous control of extremely detailed expression (RGB-level) and high-fidelity identity in portrait generation. To address the above challenges, EmojiDiff adopts a two-stage scheme involving decoupled training and fine-tuning. For decoupled training, we innovate ID-irrelevant Data Iteration (IDI) to synthesize cross-identity expression pairs by dividing and optimizing the processes of maintaining expression and altering identity, thereby ensuring stable and high-quality data generation. Training the model with this data, we effectively disentangle fine expression features in the expression template from other extraneous information (e.g., identity, skin). Subsequently, we present ID-enhanced Contrast Alignment (ICA) for further fine-tuning. ICA achieves rapid reconstruction and joint supervision of identity and expression information, thus aligning identity representations of images with and without expression control. Experimental results demonstrate that our method remarkably outperforms counterparts, achieves precise expression control with highly maintained identity, and generalizes well to various diffusion models. Code and data will be publicly available at https://emojidiff.github.io/.
@@ 1 Introduction
The development of diffusion model [20, 38, 40, 21, 32] has catalyzed significant advancements in identity (ID) customization [49, 44, 18]. This task involves generating images that align with the specified identity of portrait images, with wide-ranging applications in AI portraits [26, 29, 23], image animation [22, 46], and virtual try-ons [6, 7, 41]. Despite substantial advancement, existing methods reveal a notable limitation: monotonous facial expressions. This shortcoming constrains the expressiveness of the generated portraits, undermining their ability to convey subtle emotions and severely hindering practical applications.
@@
Researchers [10, 24, 34] seek to use reference images to manipulate portrait expressions. However, due to the strong coupling between expression and identity, it is challenging to introduce precise expression control while preserving high identity fidelity. First, fine-grained expression control conflicts with identity preservation. Excessively detailed expression templates also bring appearance-related semantic and style information (such as facial contours and skin). If the model directly replicates the appearance and structure from the expression reference, it compromises the identity of generated images (referred to as ID leakage). Second, a potential conflict exists between facial expression and identity control since both manipulate the face. Expression control can modify the facial topology and reduce identity fidelity, while identity control may make expressions toward neutrality. Third, although the highly flexible diffusion model and Style LoRA enrich the generated portraits, they also bring challenges in the generalizability of identity and expression control across various stylized scenes.
@@
Pioneering works can hardly deal with these challenges altogether. As depicted in Fig. 1, signal-conditioning generation methods [10, 24, 34] tend to extract and inject coarse expression signals (e.g., landmarks, poses) into the diffusion model to prevent ID leakage. However, these signals lose many facial details (e.g. frowning, pouting, and cheek twitching), and they are significantly limited by the accuracy of third-party networks, such as pose [2] and action unit detectors [3]. Portrait animation [22, 5] can serve as a post-processing tool for expression customization, which transfers expressions from reference templates to portrait images. However, these methods heavily rely on generated stylized images, where the identity and data distribution of the portraits significantly deviate from the real world, negatively impacting identity fidelity, expression control, and image quality.
@@
In this paper, we introduce EmojiDiff, an innovative end-to-end solution that integrates fine-grained expression control with high-fidelity portrait generation. Unlike previous signal conditioning-based methods [2, 33, 43], we still insist that the model should derive the motion directly from the original RGB expressions, ensuring comprehensive expression representation and robust generalization [17, 47]. To achieve this, we propose a pluggable expression controller (E-Adapter) alongside its decoupled training strategy. As the core of the training, we innovate ID-irrelevant Data Iteration (IDI) to synthesize the identity-expression decoupling data (i.e., cross-identity expression pairs), which are rarely available in real-world data. IDI creatively divides cross-identity expression transfer into two steps: maintaining expressions and modifying identities. By sequentially optimizing the two steps, we achieve a stable and high-quality data supply. Training the controller with these pairs facilitates the transfer of detailed expressions to the generated output while implicitly filtering out identity information concealed in the expressions, thereby preventing ID leakage. To address the potential conflict between expression and identity, we develop the ID-enhanced Contrast Alignment (ICA) for fine-tuning. ICA efficiently computes the identity and expression loss by one-step reconstruction, aligning identity representations of images across various expressions. This process mitigates the negative impacts of expression control on identity and enhances identity fidelity. In summary, our contributions are fourfold:
@@
3) We propose ID-enhanced Contrast Alignment, an efficiently fine-tuning strategy designed to ensure generated portraits maintain stable identity characteristics across diverse expressions.
4) We extensively evaluate the performance of the proposed method on different base models, styles, and expressions, showing that EmojiDiff achieves better performance compared to state-of-the-art methods.
Figure 2: Overview of the proposed method. To integrate RGB-driven expression control into diffusion models, we aim to synthesize cross-identity data $\{\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ for the model’s decoupled training, and mitigate the negative impact on the original identity through contrastive alignment fine-tuning. Before decoupled training, the fundamental expression controller (i.e., Base E-Adapter) is trained with same-identity triplet data to obtain expression transfer capabilities (the detailed structure of the E-Adapter is illustrated in Fig. 3). Next, the trained Base E-Adapter and FaceFusion [12] are utilized to alter the identity of portraits, thereby creating cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$. Subsequently, the Refined E-Adapter uses newly synthesized data for disentangled training, facilitating dual control of identity and expression without ID leakage. Finally, the Refined E-Adapter is fine-tuned by expression and identity loss based on ANI.
@@ 2.1 ID-Preserving Image Generation
The task of identity-preserving generation [40, 15, 21] seeks to empower diffusion models to synthesize images of given identities. Recent studies [26, 4, 23, 18] have explored tuning-free methods, which utilize face embeddings to conditioning generation without altering raw models. For instance, FaceChain [29] introduces an independent adapter before the cross-attention layer, preventing interference between face and text conditions. IP-Adapter [49] introduces a novel decoupled cross-attention mechanism for identity injection. InstantID [44] proposes an IdentityNet to incorporate facial key points with image conditions. Although these methods can quickly generate customized portraits, they severely lack fine control over the expressions.
@@ 2.2 Expression Customization
Most methods for expression animation integrate control signals extracted from expression images into diffusion models. ControlNet [51] is a pioneering effort that enables the incorporation of human poses, depth maps, and other image conditions. Following this framework, MediaPipe [8] specializes in extracting facial landmarks from reference expressions to preserve more expression details. Arc2Face [10] extracts 3D normal maps [14] to personalize expressions based on given portraits. Additionally, some studies pursue more flexible control. Collaborative Diffusion [24] allows localized editing of facial parts using semantic maps. [34] combined valence & arousal, and action units and GANmut [9] to characterize emotion information. DiffSFSR [28] utilizes an emotional dictionary to transfer facial expressions to the generated image. FineFace [43] employs action units to allow localized control by describing the intensity of facial muscle movements.
In practice, portrait animation [50, 22, 31, 11, 5, 45, 46] can adjust portrait expressions without altering the image background. Notable examples, such as FaceAdapter [19], X-Portrait [47], and LivePortrait [17], which can be used for post-processing of generated portrait images. However, these methods cannot simultaneously customize real-world identities and expressions, adversely affecting identity fidelity and image quality.
@@ 3 Methodology
Given a source portrait $\mathbf{S}$, a textual description $\mathbf{T}$, and a reference $\mathbf{R}$, our method aims to generate customized images $\mathbf{D}$. To enable the model to accept RGB expression inputs, we aim to construct high-quality ternary data sets $\mathcal{T}=\{\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ for decoupled training, where $k$ and $g$ denote the $k$-th ID and the $g$-th expression, respectively. During training, the model takes $\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g}$ as inputs and $\mathbf{D}_{k}^{g}$ as the learning objective, represented as $\mathbf{D}_{k}^{g}=\mathcal{F}(\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k^{\prime}}^{g})$, where $\mathcal{F}$ is the trained expression controller. This setup allows the model to maintain the identity of $\mathbf{S}$ while accurately replicating the expression from $\mathbf{R}$. However, we are unable to collect qualified cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ from real-world scenarios. Besides, the integration of expression control hinders the original ID control.
To overcome these issues, our method is divided into two stages: decoupling training and fine-tuning. Before decoupled training, we construct the required data smartly in two steps. Firstly, we design the expression controller and train a base version (Base E-Adapter) using same-identity ternary data sets as detailed in Sec. 3.2.1. The images generated by Base E-Adapter exhibit a very high expression consistency with given expression references, although they also have a similar appearance to the references (ID leakage). Secondly, we develop the ID-irrelevant Data Iteration (IDI) to alter the identity of the images generated by Base E-Adapter, thus promoting the given reference expressions and the altered images to form data pairs with the same expression but different identity (as outlined in Sec. 3.2.2). Using the newly created cross-identity dataset, we perform decoupled training on the proposed expression controller to develop the Refined E-Adapter. Refined E-Adapter accurately controls expression details while preventing ID leakage when generating portrait images (see Sec. 3.2.3). Finally, we introduce ID-enhanced Contrast Alignment (ICA) to further fine-tune the Refined E-Adapter, reducing the negative impact of expression manipulation on ID control (Sec. 3.3). Fig. 2 provides an overview of our pipeline.
@@ 3.1 Preliminary
Latent Diffusion Model. The latent diffusion models involve a diffusion process and a reverse process in the latent space. During the diffusion process, an image $x$ is first projected to a smaller latent representation by a VAE [25] encoder. Then, random noise is gradually added to the representation $z_{0}$. The noisy representation $z_{t}$ is derived as:
@@
where $t$, $\bar{\alpha}_{t}$ denotes timestep and predefined function of $t$, respectively. Conversely, during the denoising process, a U-Net [39] is trained to predict the added noise $\epsilon$ from the noisy latent representation $z_{t}$. The training objective aims to minimize the loss function as follows:
@@
Image Prompt Adapter. IP-Adapter [49] introduces a novel approach to incorporating an image prompt, working together with a text prompt to control image generation. A decoupled cross-attention is used to separate cross-attention layers for text features and image features, formulated as:
@@ 3.2 Decoupled Training
Cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ are essential for training expression controllers. However, we cannot obtain adequate samples from the real world due to the subtle differences in individual expressions. Additionally, existing methods [17, 12, 47, 48] also struggle to synthesize qualified data pairs, particularly with extreme expressions. Therefore, we seek to develop a novel strategy for data construction.  Reviewing previous works [47, 48], we notice that they avoid using same-identity expression pairs $\{\mathbf{R}_{k}^{g},\mathbf{D}_{k}^{g}\}$ for controller training to prevent identity leakage during inference. However, our insights suggest that the controller trained with same-identity pairs exhibits superior expression transfer capabilities, as shown in Fig. 4. Based upon this, we divide the synthesis of cross-identity expression pairs into two sequential steps: maintaining expression and modifying identity. Initially, we utilize same-identity data to train an expression controller that accurately transfers reference expressions into generated images without considering identity leakage. Subsequently, we apply portrait generation techniques to further alter the identity of the controller-generated image. With the powerful abilities of the diffusion model, the modified image and the given expression template form a high-quality cross-identity data pair. After that, the data can be used to train the model to decouple expressions from identities.

### v3

@@ Abstract
This paper aims to bring fine-grained expression control while maintaining high-fidelity identity in portrait generation. This is challenging due to the mutual interference between expression and identity. On one hand, fine expression control signals inevitably introduce appearance-related semantics (e.g., facial contours, and ratio), which impact the identity of the generated portrait. On the other hand, even coarse-grained expression control can cause facial changes that compromise identity, since they all act on the face. Here, we introduce EmojiDiff, the first end-to-end solution that enables simultaneous control of extremely detailed expression (RGB-level) and high-fidelity identity in portrait generation. To address the above challenges, EmojiDiff adopts a two-stage scheme involving decoupled training and fine-tuning. For decoupled training, we innovate ID-irrelevant Data Iteration (IDI) to synthesize high-quality cross-identity expression pairs by separating and optimizing the processes of maintaining expression and altering identity. Training the model with this data, we effectively disentangle fine expression features in the expression template from other extraneous information (e.g., identity, skin). Subsequently, we present ID-enhanced Contrast Alignment (ICA) for further fine-tuning. ICA achieves rapid reconstruction and joint supervision of identity and expression information, thus aligning identity representations of images with and without expression control. Experimental results demonstrate that our method significantly outperforms its counterparts, achieving precise expression control with highly maintained identity, and generalizing well to various diffusion models. Project page: https://emojidiff.github.io/.
† Corresponding authors.
@@ 1 Introduction
The development of diffusion model [ho2020denoising, rombach2022sd, ruiz2023dreambooth, hu2021lora, nichol2021glide] has catalyzed significant advancements in identity (ID) customization [ye2023ipa, wang2024instantid, guo2024pulid]. This task involves generating images that align with the specified identity of portrait images, with wide-ranging applications in AI portraits [li2024photomaker, liu2023facechain, huang2024consistentid], image animation [hu2024animateanyone, wei2024aniportrait], and virtual try-ons [choi2024vton, chong2024catvton, shen2024imagdressing]. Despite substantial advancement, existing methods reveal a notable limitation: monotonous facial expressions. This shortcoming constrains the expressiveness of the generated portraits, undermining their ability to convey subtle emotions and severely hindering practical applications.
@@
Researchers [deng2019arcface, huang2023collaborative, paskaleva2024unified] seek to use reference images to manipulate portrait expressions. However, due to the strong coupling between expression and identity, it is challenging to introduce precise expression control while preserving high identity fidelity. First, fine-grained expression control conflicts with identity preservation. Excessively detailed expression templates also bring appearance-related semantic and style information (such as facial contours and skin). If the model directly replicates the appearance and structure from the expression reference, it compromises the identity of generated images (referred to as ID leakage). Second, a potential conflict exists between facial expression and identity control since both manipulate the face. Expression control can modify the facial topology and reduce identity fidelity, while identity control may make expressions more neutral. Third, although the highly flexible diffusion model and Style LoRA enrich the generated portraits, they also bring challenges in the generalizability of identity and expression control across various stylized scenes.
@@
Pioneering works can hardly deal with these challenges altogether. As depicted in Fig. 1, signal-conditioning generation methods [deng2019arcface, huang2023collaborative, paskaleva2024unified] tend to extract and inject coarse expression signals (e.g., landmarks, poses) into the diffusion model to prevent ID leakage. However, these signals lose many facial details (e.g., frowning, pouting, and cheek twitching), and they are significantly limited by the accuracy of third-party networks, such as pose [OpenPose] and action unit detectors [chang2023libreface]. Portrait animation [hu2024animateanyone, chen2024echomimic] can serve as a post-processing tool for expression customization, which transfers expressions from reference templates to portrait images. However, these methods heavily rely on generated stylized images, where the identity and data distribution of the portraits significantly deviate from the real world, negatively impacting identity fidelity, expression control, and image quality.
@@
In this paper, we introduce EmojiDiff, an innovative end-to-end solution that integrates fine-grained expression control with high-fidelity portrait generation. Unlike previous signal conditioning-based methods [OpenPose, paraperas2024arc2face, varanka2024fineface], we still insist that the model should derive the motion directly from the original RGB expressions, ensuring comprehensive expression representation and robust generalization [guo2024liveportrait, xie2024xportrait]. To achieve this, we propose a pluggable expression controller (E-Adapter) alongside its decoupled training strategy. As the core of the training, we innovate ID-irrelevant Data Iteration (IDI) to synthesize the identity-expression decoupling data (i.e., cross-identity expression pairs), which are rarely available in real-world data. IDI creatively divides cross-identity expression transfer into two steps: maintaining expressions and modifying identities. By sequentially optimizing the two steps, we achieve a stable and high-quality data supply. Training the controller with these pairs facilitates the transfer of detailed expressions to the generated output while implicitly filtering out identity information concealed in the expressions, thereby preventing ID leakage. To address the potential conflict between expression and identity, we develop the ID-enhanced Contrast Alignment (ICA) for fine-tuning. ICA efficiently computes the identity and expression loss by one-step reconstruction, promoting the identity under different expressions as consistent as possible. This process mitigates the negative impacts of expression control on identity and enhances identity fidelity. In summary, our contributions are fourfold:
@@
3) We propose ID-enhanced Contrast Alignment, an efficient fine-tuning strategy designed to ensure generated portraits maintain stable identity across diverse expressions.
4) We introduce the cross-identity expression dataset CIEP100k, containing various expressions and data with the same expression but different identities, facilitating further research on expression-identity disentanglement.
Figure 2: Overview of the proposed method. To integrate RGB-driven expression control into diffusion models, we aim to synthesize cross-identity data $\{\mathbf{S}_{i}^{\neg{e}},\mathbf{R}_{\neg{i}}^{e},\mathbf{T}_{i}^{e}\}$ for the model’s decoupled training, and mitigate the negative impact on the original identity through contrastive alignment fine-tuning. Before decoupled training, the fundamental expression controller (i.e., Base E-Adapter) is trained with same-identity data $\{\mathbf{S}_{i}^{\neg{e}},\mathbf{R}_{i}^{e},\mathbf{T}_{i}^{e}\}$ to obtain expression transfer capabilities (the structure of the E-Adapter is illustrated in Fig. 3). Next, the trained Base E-Adapter and FaceFusion [FaceFusion] are utilized to alter the identity of portraits, thereby creating cross-identity expression pairs $\{\mathbf{R}_{\neg{i}}^{e},\mathbf{T}_{i}^{e}\}$. Subsequently, the Refined E-Adapter uses newly synthesized data for disentangled training, facilitating dual control of identity and expression without ID leakage. Finally, the Refined E-Adapter is fine-tuned by expression and identity loss based on ANI.
@@ 2.1 ID-Preserving Image Generation
The task of identity-preserving generation [ruiz2023dreambooth, gal2022inversion, hu2021lora] seeks to empower diffusion models to synthesize images of given identities. Recent studies [li2024photomaker, chen2023photoverse, huang2024consistentid, guo2024pulid] have explored tuning-free methods, which utilize face embeddings to conditioning generation without altering raw models. For instance, FaceChain [liu2023facechain] introduces an independent adapter before the cross-attention layer, preventing interference between face and text conditions. IP-Adapter [ye2023ipa] introduces a novel decoupled cross-attention mechanism for identity injection. InstantID [wang2024instantid] proposes an IdentityNet to incorporate facial key points with image conditions. Although these methods can quickly generate customized portraits, they severely lack fine control over the expressions.
@@ 2.2 Expression Customization
Most methods for expression animation integrate control signals extracted from expression images into diffusion models. ControlNet [zhang2023controlnet] is a pioneering effort that enables the incorporation of human poses, depth maps, and other image conditions. Following this framework, MediaPipe [MediaPipeFace] specializes in extracting facial landmarks from reference expressions to preserve more expression details. Arc2Face [deng2019arcface] extracts 3D normal maps [filntisis2022emoca] to personalize expressions based on given portraits. Additionally, some studies pursue more flexible control. Collaborative Diffusion [huang2023collaborative] allows localized editing of facial parts using semantic maps. [paskaleva2024unified] combined valence & arousal, and action units and GANmut [d2021ganmut] to characterize emotion information. DiffSFSR [liu2024DiffSFSR] utilizes an emotional dictionary to transfer facial expressions to the generated image. FineFace [varanka2024fineface] employs action units to allow localized control by describing the intensity of facial muscle movements.
In practice, portrait animation [zeng2023fadm, hu2024animateanyone, ma2024followyouremoji, drobyshev2024emoportraits, chen2024echomimic, wang2024instructavatar, wei2024aniportrait] can adjust portrait expressions without altering the image background. Notable examples, such as FaceAdapter [han2024faceadapter], X-Portrait [xie2024xportrait], and LivePortrait [guo2024liveportrait], which can be used for post-processing of generated portrait images. However, these methods cannot simultaneously customize real-world identities and expressions, adversely affecting identity fidelity and image quality.
@@ 3 Methodology
Given a source portrait $\mathbf{S}$, a textual prompt $\mathbf{P}$, and a reference $\mathbf{R}$, our method aims to generate target images $\mathbf{T}$. To enable the model to accept RGB expression inputs, we aim to construct ternary data sets $\mathcal{T}=\{\mathbf{S}_{i}^{\neg{e}},\mathbf{R}_{\neg{i}}^{e},\mathbf{T}_{i}^{e}\}$ for decoupled training. Here, $i$ and $e$ denote the specific identity and the expression, and $\neg{i}$ and $\neg{e}$ can be any specific items different from $i$ and $e$. During training, the model takes $\mathbf{S}_{i}^{\neg{e}},\mathbf{R}_{\neg{i}}^{e}$ as inputs and $\mathbf{T}_{i}^{e}$ as the learning objective, represented as $\mathbf{T}_{i}^{e}=\mathcal{F}(\mathbf{S}_{i}^{\neg{e}},\mathbf{R}_{\neg{i}}^{e})$, where $\mathcal{F}$ is the trained expression controller. This setup allows the model to maintain the identity $i$ of $\mathbf{S}$ while accurately replicating the expression $e$ from $\mathbf{R}$. However, we are unable to collect qualified cross-identity expression pairs $\{\mathbf{R}_{\neg{i}}^{e},\mathbf{T}_{i}^{e}\}$ from real-world scenarios. Besides, the integration of expression control hinders the original ID control.
To overcome these issues, our method is divided into two stages: decoupling training and fine-tuning. Before decoupled training, we construct the required data delicately in two steps. Firstly, we design the expression controller and train a base version (Base E-Adapter) using the same-identity ternary data sets as detailed in Sec. 3.2.1. The images generated by Base E-Adapter exhibit a very high expression consistency with given expression references, although they also have a similar appearance to the references (ID leakage). Secondly, we develop the ID-irrelevant Data Iteration (IDI) to alter the identity of the images generated by Base E-Adapter, thus promoting the given reference expressions and the altered images to form data pairs with the same expression but different identity (as outlined in Sec. 3.2.2). Using the newly created cross-identity dataset, we perform decoupled training on the proposed expression controller to develop the Refined E-Adapter. Refined E-Adapter accurately controls expression details while preventing ID leakage when generating portrait images (see Sec. 3.2.3). Finally, we introduce ID-enhanced Contrast Alignment (ICA) to further fine-tune the Refined E-Adapter, reducing the negative impact of expression manipulation on ID control (Sec. 3.3). Fig. 2 provides an overview of our pipeline.
@@ 3.1 Preliminary
Latent Diffusion Model. The latent diffusion models involve a diffusion process and a reverse process in the latent space. During the diffusion process, an image $x$ is first projected to a smaller latent representation by a VAE [kingma2013vae] encoder. Then, random noise is gradually added to the representation $z_{0}$. The noisy representation $z_{t}$ is derived as:
@@
where $t$, $\bar{\alpha}_{t}$ denotes timestep and predefined function of $t$, respectively. Conversely, during the denoising process, a U-Net [ronneberger2015unet] is trained to predict the added noise $\epsilon$ from the noisy latent representation $z_{t}$. The training objective aims to minimize the loss function as follows:
@@
Image Prompt Adapter. IP-Adapter [ye2023ipa] introduces a novel approach to incorporating an image prompt, working together with a text prompt to control image generation. A decoupled cross-attention is used to separate cross-attention layers for text features and image features, formulated as:
@@ 3.2 Decoupled Training
Cross-identity expression pairs $\{\mathbf{R}_{\neg{i}}^{e},\mathbf{T}_{i}^{e}\}$ are essential for training expression controllers. However, we cannot obtain adequate samples from the real world due to the subtle differences in individual expressions. Additionally, existing methods [guo2024liveportrait, FaceFusion, xie2024xportrait, yang2024megactor] also struggle to synthesize qualified data pairs, particularly with extreme expressions. Therefore, we seek to develop a novel strategy for data construction.  Reviewing previous works [xie2024xportrait, yang2024megactor], we notice that they avoid using same-identity expression pairs $\{\mathbf{R}_{i}^{e},\mathbf{T}_{i}^{e}\}$ for controller training to prevent identity leakage during inference. However, our insights suggest that the controller trained with same-identity pairs exhibits superior expression transfer capabilities, as shown in Fig. 4. Based upon this, we divide the synthesis of cross-identity expression pairs into two sequential steps: maintaining expression and modifying identity. Initially, we utilize the same-identity data to train an expression controller that accurately transfers reference expressions into generated images without considering identity leakage. Subsequently, we utilize the portrait generation process to further alter the identity of the controller-generated image. With the powerful abilities of the diffusion model, the modified image and the given expression template form a high-quality cross-identity data pair. After that, the data can be used to train the model to decouple expressions from identities.